# Scaling Laws and Compute-Optimal Training Beyond Fixed Training Durations

### v2

@@ 2 Background: Cosine Learning Rate Schedule for LLMs
Experimental visualization. Our first goal is to understand the importance of the length of the schedule for performance of the model. To this end, we implement the common decoder-only transformer (Vaswani et al., 2017) identical to the LLaMa (Touvron et al., 2023a, b) or Noam architecture (Ormazabal et al., 2024). Throughout this paper, we use the AdamW optimizer with weight decay (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with common LLM training parameters. We train on a subset of SlimPajama (Soboleva et al., 2023) with 6B tokens [^1] https://huggingface.co/datasets/DKYoon/SlimPajama-6B, a cleaned and deduplicated corpus for LLM pretraining, which we split into train and validation sequences and report validation loss (perplexity). We provide all details in Appendix A.1.
@@ 3 A Different Route: Constant Learning Rate with Cooldown
Figure 2: Illustration of schedules. Cosine (red) follows a slow decrease in learning rate, typically to 10% of the maximum for LLMs. The alternative (blue) is characterized by an aggressive decrease in learning rate via e.g. a linear cooldown.
@@
The alternative — constant + cooldown. The tradeoff of cosine can also be achieved by a different schedule. Here, the learning rate is kept constant for the majority of the training and only decreases in a short final phase, known as the cooldown (or decay/annealing phase). This schedule has previously been referred to as a trapezoidal (Zhai et al., 2022), and later as the warmup-stable-decay schedule (WSD) (Hu et al., 2024). To avoid overloading of terms (e.g., weight decay), we refer to this approach as constant LR + cooldown.
The cooldown phase typically has the LR go to zero linearly, mirroring the warmup phase, which gives rise to an overall trapezoidal shape (see Figure 2). Formally, we can define it as:
@@
Conceptual advantages. The main advantage of this schedule is that specifying the number of training steps in advance is not required. This is particularly convenient for large runs, as the cooldown can be initiated at any time to observe model behavior and decide whether to stop. It also allows for continual learning by default, for which training can be resumed from a checkpoint prior cooldown. Moreover, the data mixture can be changed during the cooldown phase (Hu et al., 2024) as a form of finetuning; while we focus on the same mixture, understanding the curriculum aspect of the separate phases is an important direction for future research.
Experimental comparison. We follow our experimental setup and train a 210M parameter model on SlimPajama with constant LR and the cooldown schedule defined in (1). That is, we compare the same length of warmup and training, but replace the cosine decay after warmup with a constant LR and a cooldown for 20% of steps linearly going to zero. We additionally sweep the learning rate for both approaches. In the results shown in Figure 3, perhaps suprisingly, we observe an almost perfect match between the performance of the best cosine and cooldown schedule even for different training durations, all while exhibiting slightly less sensitivity to variations in the LR.
Different cooldown schedules. We investigate various functions describing the cooldown shape, including cosine, square and square root shapes in Appendix B.1. We identify a new function (1-sqrt) that outperforms the linear cooldown. This improvement is maintained in a smaller number of decay steps, different learning rates, and various timestamps (see Appendix B.1). In Figure 4, we train a model for 200K steps (approximately 20B tokens), applying cooldown every 20K steps for 20%. The results indicate that the longer the training duration, the more the linear cooldown is outperformed by the (1-sqrt) cooldown, which we define as:
@@
Takeaway 1: The constant LR + cooldown schedule offers significant convenience by not requiring the number of training steps to be specified in advance, and provides similar or better performance compared to the optimally tuned cosine schedule.
@@
Takeaway 2: We introduce a cooldown form (1-sqrt) that consistently outperforms the linear decay.
@@
Figure 4: A different cooldown schedule can improve performance. Perhaps surprisingly, we find that a different decay phase in the functional form of (1-sqrt) can consistently outperforms the standard linear decay, where both are better than a well-tuned cosine for long lengths.
@@
How long do you need to cooldown? To effectively utilize the schedule, it is essential to determine the optimal number of decay steps. Our study on the relative number of cooldown steps, as shown in Fig. 5, reveals that the benefits of extended cooldown periods plateau at around 20%. We select this percentage for our experiments as the model size and the number of tokens are relatively small. Additionally, in Fig. 6 we demonstrate that using only 5% decay with the (1-sqrt) cooldown can nearly match the performance of the cosine schedule on a 20B token run (much beyond Chinchilla optimal).
@@
Figure 5: Longer cooldown helps to achieve lower loss. We investigate the effect of the cooldown length as a fraction of the total steps for both the 210M model (left) and 60M (right). For a well-tuned learning rate of cosine, we find that the cooldown surpasses cosine between 10-20% of steps (left), but stops improving when done over a majority of training. This also holds when sweeping the learning rate (right). Additional ablations are provided in Appendix B.1.
@@
Figure 6: A long training run suggests that a small number of cooldown steps can match cosine for long training. From Fig. 5 and Fig. 17, we find that the required duration of cooldown to match the cosine loss decreases with longer training; we validate with a long training run (200k steps), and find that just 10k cooldown steps almost perfectly match cosine in performance.
@@
Takeaway 3: The constant learning rate with a short cooldown (< 20% of total training steps) achieves the same final loss as a well-tuned cosine schedule, and only outperforms for a longer fraction of cooldown steps. For long training runs, our results suggest that the cooldown length can be less than 20% to match cosine if enough in absolute steps — see Figure 6.
@@
What happens during the cooldown? It is remarkable how the sudden drop in loss is consistent for both train and validation loss and aligns closely with the decay in learning rate (Figure 13). Hu et al. (2024) investigate the cooldown phase and find that the first-order directional derivative diminishes with each step, whereas the curvature of the loss function increases; they attribute this to proximity to a local optimum.
@@
We expand upon these results and aim to understand the optimization landscape around the trajectory of the cooldown. For that, we evaluate the loss along the straight line trajectory when moving between a checkpoint before and after cooldown, i.e., a linear interpolation of the two models’ weights. Perhaps surprisingly, we find that the smooth drop in loss also occurs for this interpolation, as visualized in Figure 7. This aligns with the findings of Hu et al. (2024). These results suggest that upon decaying the learning rate, the model immediately descends into a connected minimum of the loss.
@@ 4 Do We Even Need to Cooldown?
In Section 3, we showed that a constant LR with a short cooldown phase can replace the cosine decay. Ideally, however, we would not need a decay at all, but aim to obtain an optimal model at any point in training even when prolonged. This would save even more computational resources and time. In this section, we investigate two potential approaches: weight averaging and a schedule-free optimizer.
@@ 4.1 Stochastic Weight Averaging (SWA)
Motivation. While slow annealing of the learning rate can be essential to find good minima (Smith et al., 2017; Loshchilov & Hutter, 2016), Sandler et al. (2023) show theoretical and empirical equivalence between stochastic weight averaging (SWA) and cooldown schedules in training vision models. There, averaging intuivitely and naturally reduces noise and thereby improves generalization. Motivated by these results, we aim to answer the same question in LLM training: Can weight averaging replace the cooldown phase?
Method. We opt for a form of SWA (Izmailov et al., 2018) that splits the training in fixed windows and averages within a window, which allows to keep the average as a single additional copy of the model parameters. In our experiments, we set the window to $h=500$ steps and save checkpoints every $h$ steps, which allows evaluating longer windows ad-hoc to see potential improvements akin to latest weight averaging (LAWA, Kaddour, 2022; Sanyal et al., 2023). For all our experiments, we find that windows below or at 2500 steps (256M tokens) are optimal. We also experimented with an exponential moving average (EMA), which performed worse than SWA, and therefore do not report EMA.
Experimental results. We evaluate SWA in the same setup of the 210M model and show the results in Fig. 8 for both a constant LR (left) and cosine (right). Notably, we find a significant performance boost for SWA on top of a constant LR. Yet, it does not reach the loss values of cooldowns at intermediate steps. On the other hand, in line with previous work (Kaddour, 2022; Sanyal et al., 2023), we see a similar boost for SWA on top of a cosine schedule. This suggests that SWA, regardless of schedule, provides a compelling approach to achieving strong models along the data-scale axis that can serve as a replacement for models trained with less steps, if the performance gap is acceptable and one wishes to avoid cooldowns. This is particularly advantageous as it can arguably be done for free on top of existing and well-tuned optimizers.
@@ 4.2 Schedule-Free Optimizer (SFO)
Results. We compare the results of a long 210M model run with cooldown vs. SFO with AdamW in Figure 9. While the optimizer does not require a learning rate schedule, the authors point out that training is more sensitive to the choice of the $(\beta_{1},\beta_{2})$ momentum parameters (which are not identical to the momentum in Adam). They note that the optimal parameters may depend on the length of training, making it not fully schedule free. We observe this sensitivity in our experiments, where the choice of $(0.9,0.95)$ performs notably worse and even increases loss towards the end of training. For $(\beta_{1}=0.95,\beta_{2}=0.99)$, SFO performs remarkably well. Nonetheless, both settings are matched or outperformed by the cooldown schedule, in particular when comparing the same configuration of momentum. We did not perform further hyperparameter tuning for either method.
@@ 5 The Implications for Scaling Law Research
where $\{A,\alpha,B,\beta,E\}$ are variables to be estimated (Kaplan et al., 2020; Hoffmann et al., 2022). Such scaling laws serve a multitude of critical purposes — from optimally spending a fixed amount of compute (FLOPs) for achieving the lowest loss, to trading off data-sources of different quality (Bi et al., 2024; Goyal et al., 2024) or modalities (Aghajanyan et al., 2023), to comparing the efficiency of different architectures (Gu & Dao, 2023; De et al., 2024).
Crucially, Hoffmann et al. (2022) demonstrated that is necessary to vary the number of training steps (tokens) for a fixed family of models. At the time, their results suggested that LLMs were over-sized, leading to a substantial increase in data collection and training for much longer, beyond the Chinchilla optimal point $(N,D)$ (Touvron et al., 2023a, b).
@@
Experimental setup. We mimic a small-scale experimental setup for scaling laws: We train a range of model sizes (33M-360M) across different token scales (0.3B-10B) on the same SlimPajama 6B dataset. For each model, we choose exactly three token counts (around the Chinchilla optimal ratio of D/N=20) in increments of 10, 20 and 30 tokens per parameter. With the cosine schedule, each model is trained from scratch three times. In contrast, for averaging, we train each model just once for the longest cycle, then use the averages at the same token count as cosine; for cooldown, we similarly take checkpoints along the constant LR trajectory and perform three annealing periods to match the token counts. We adjust the LR to be higher for smaller models. In line with the findings of Sect. 3 and for a fair comparison, we set the constant LR to be half of the maximum LR for cosine and perform 20% (linear) cooldown steps as the longest runs are below 70k steps. More details are given in Appendix A.1.
@@
Figure 10: Cooldown LR schedule + SWA scale reliably. To validate our findings, we launch a range of model sizes (33M-360M), each for three different cosine cycle lengths. For the other methods, we train each model just once for the longest cycle and take the snapshots at the same token count as cosine (for SWA) or perform post-train cooldowns to the same length. Each final models is represented by a dot. Left: The loss curve envelopes. Right: The perplexity of cosine (y-axis) vs. cooldowns and SWA. Points on the diagonal indicate the same loss for both methods; above outperforming cosine, below fall short. We see alignment between both methods.
@@
Additional results. We plot the training curves of all models in Appx. B.2. In addition, we show that our findings transfer to different datasets with experiments on OpenWebText2 in Appx. B.4.

### v3

@@ 2 Background: Cosine Learning Rate Schedule for LLMs
Experimental setup. Our first goal is to understand the importance of the length of the schedule for performance of the model. To this end, we implement the common decoder-only transformer (Vaswani et al., 2017) identical to the LLaMa (Touvron et al., 2023a, b) or Noam architecture (Ormazabal et al., 2024). Throughout this paper, we use the AdamW optimizer with weight decay (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with common LLM training parameters. We train on a subset of SlimPajama (Soboleva et al., 2023) with 6B tokens [^1] https://huggingface.co/datasets/DKYoon/SlimPajama-6B, a cleaned and deduplicated corpus for LLM pretraining, which we split into train and validation sequences and report validation loss (perplexity). We provide all the details in Appendix A.1.
@@ 3 A Different Route: Constant Learning Rate with Cooldown
Figure 2: Illustration of schedules. Cosine (red) follows a slow decrease in learning rate, typically to 10% of the maximum for LLMs. The alternative is characterized by an aggressive decrease in learning rate, e.g., via a linear (blue) or square root (yellow) cooldown.
@@
The alternative — constant + cooldown. The tradeoff of cosine can also be achieved by a different schedule. Here, the learning rate (LR) is kept constant for the majority of the training and only decreases in a short final phase, known as the cooldown (or decay/annealing phase). This schedule has previously been referred to as a trapezoidal (Zhai et al., 2022), and later as the warmup-stable-decay schedule (WSD) (Hu et al., 2024). To avoid overloading of terms (e.g., weight decay), we refer to this approach as constant LR + cooldown.
The cooldown phase typically has the LR go to zero, mirroring the warmup phase, which gives rise to an overall trapezoidal shape (see Figure 2). Formally, we can define it as
@@
3.1 The Advantages
The main advantage of the constant schedule is that it does not require one to specify the number of training steps in advance. This is particularly convenient for large runs, as the cooldown can be initiated at any time to observe model behavior and decide whether to stop. We want to highlight the following in more detail.
Continual learning. The constant + cooldown schedule allows for continual learning by default. Here, the natural approach is to use checkpoints before the cooldown to continue training with a high LR; this avoids loss spikes and brittle training when rewarming the learning rate (cf. Figure 1 with cosine). Moreover, rewarming has been reported to hurt performance and introduce forgetting compared to single annealing training (Ibrahim et al., 2024), though careful strategies can alleviate such issues. It remains an interesting question if a single cooldown schedule is absolutely optimal given a total compute budget for LLM training.
Data mixtures. During the cooling phase, the data mixture can be changed (Hu et al., 2024) as a form of finetuning; beyond aiming for specific downstream tasks, this allows one to assess the quality of specific mixes, e.g., as recently done for Llama 3 (Dubey et al., 2024). Although we focus on the same mixture, understanding the curriculum aspect of the separate phases is explored in the concurrent work of Blakeney et al. (2024) and remains open for further research.
Scaling studies. Since a cooldown can also be initiated retrospectively from a checkpoint, the schedule allows to see the same model at different scales of compute. This enables much cheaper scaling studies than commonly done; we show this in depth in Section 5.
3.2 Experimental Comparison
We follow our experimental setup from the previous section (details in A.1) and train a 210M parameter model on SlimPajama with constant LR and the cooldown schedule defined in (1). That is, we compare the same length of warmup and training, but replace the cosine decay after warmup with a constant LR and a cooldown for 20% of steps linearly going to zero. We additionally sweep the maximum learning rate for both approaches. In the results shown in Figure 3, perhaps suprisingly, we observe an almost perfect match between the performance of the best cosine and cooldown schedule even for different training durations, all while exhibiting slightly less sensitivity to variations in the LR.
Different cooldown schedules. We investigate various functions describing the cooldown shape, including cosine, square, and square root shapes in Appendix B.1. We identify a new function (1-sqrt) that outperforms the linear cooldown. This improvement is maintained in a smaller number of decay steps, different learning rates, and various timestamps (see Appendix B.1). In Figure 4, we train a model for 200K steps (approximately 20B tokens), applying cooldown every 20K steps for 20%. The results indicate that the longer the training duration, the more linear cooldown is outperformed by the (1-sqrt) cooldown, which we define as:
@@ 3.2 Experimental Comparison
Takeaway 1: The constant LR + cooldown schedule offers significant convenience by not requiring the number of training steps to be specified in advance, and provides similar performance compared to a well tuned cosine schedule.
@@
Takeaway 2: We find a cooldown form (1-sqrt) that consistently performs better than linear decay.
@@
Figure 4: A different cooldown schedule can improve performance. Perhaps surprisingly, we find that a different decay phase in the functional form of (1-sqrt) can consistently outperform the standard linear decay, where both are better than the chosen (untuned) cosine for long lengths.
@@
How long do you need to cooldown? To effectively utilize the schedule, it is essential to determine the optimal number of decay steps. Our study of the relative number of cooldown steps, as shown in Fig. 5, reveals that the benefits of extended cooldown periods plateau at around 20%, which we select for the remaining experiments. Additionally, in Fig. 6 we demonstrate that using only 5% decay with the (1-sqrt) cooldown can nearly match the performance of the cosine schedule on a 20B token run (much beyond Chinchilla optimal). This is practically important, as we ideally keep the cooldown as short as possible but get strong performance.
@@
Figure 5: Longer cooldown helps to achieve lower loss. We investigate the effect of the cooldown length as a fraction of the total steps for a 124M model. We find that the cooldown surpasses cosine between 10-20% of steps (left), but largely stops improving when done over a majority of training. This also holds when sweeping the LR (right). Additional ablations are provided in Appendix B.1.
@@
Figure 6: A long training run suggests that a small number of cooldown steps can match cosine for long training. From Fig. 5 and Fig. 20, we find that the required duration of cooldown to match the cosine loss decreases with longer training; we validate with a long training run (200k steps, same LRs), and find that just 10k cooldown steps almost perfectly match cosine in performance.
@@
Takeaway 3: The constant learning rate with a short cooldown (< 20% of total training steps) can achieve the same final loss as cosine, and only outperforms for a longer fraction of cooldown steps. For long training runs, our results suggest that the cooldown length can be less than 20% to match cosine if it is enough in absolute steps — see Figure 6.
@@
What happens during the cooldown? It is remarkable how the sudden drop in loss is consistent for both train and validation loss and aligns closely with the decay in LR (Figure 15). Hu et al. (2024) investigate the cooldown phase and find that the first-order directional derivative diminishes with each step, whereas the curvature of the loss function increases; they attribute this to proximity to a local optimum.
@@
We expand upon these results and aim to understand the optimization landscape around the trajectory of the cooldown. For that, we evaluate the loss along the straight line trajectory when moving between a checkpoint before and after cooldown, i.e., a linear interpolation of the weights of the two models. Perhaps surprisingly, we find that the smooth drop in loss also occurs for this interpolation, as visualized in Figure 7. This aligns with the findings of Hu et al. (2024). These results suggest that upon decaying the LR, the model immediately descends into a connected minimum of the loss.
@@
Decaying cosine to less than 10%. Throughout our experiments, we follow the common heuristic of setting the final LR of cosine to 10% of the maximum. We also provide experiments in which we ablate annealing to zero (or a very small value) in Fig. 22 in Appendix B.1. We find that, perhaps unsurprisingly, the influence of the final LR is non-negligible: decaying to lower values improves performance to match the best (1-sqrt) cooldown. However, when evaluating on downstream benchmarks (see next Section 3.3), we see that annealing to zero can hurt metrics; we posit this comes from too early saturation. Combining these two arguments implies that the maximum and final LR should be set (and ideally swept over) independently.
(a) Metrics throughout training.
Aggregated Score
Cosine to 10% 46.26
Cosine to 0 45.88
Linear (20%) 46.20
(1-Sqrt) (20%) 46.23
(a) Final scores (100B tokens).
Figure 8: The performances of both schedules also match for downstream tasks. We run a realistic setting of a 1B model trained for 100B tokens of FineWeb and establish matching performance of both schedules for common LLM benchmarks. Interestingly, there is a similar boost in performance with the cooldown. Detailed numbers over the course of training and 460B token runs are in Appx. B.5.
3.3 Scaling Up: 1B and 8B Models
We expand our ablations to realistic scenarios of 1B and 8B parameter models, with an additional focus on model evaluations. Our goal in this section is therefore twofold: first, we validate the dynamics of the schedule at practical scale; second, we establish the connection between the pure loss values and downstream benchmarks. Although tasks often scale reliably with loss (Du et al., 2024; Gadre et al., 2024), this connection is crucial, as downstream abilities are ultimately the main metric of interest.
[…]
Figure 9: Results at 8B scale. We validate the behavior with a much larger model (architecture of Llama 3) for a single short run (12B tokens of FineWeb-Edu), where the cooldown matches the cosine schedule again.
With the aggregated results in Fig. 8, we find a clear agreement of the downstream performance for both schedules. Interestingly, there is an uptick for some evaluations, similar to the drop in loss, with the start of the LR cooldown; see Fig. 29. However, not all benchmarks follow this trend, which requires further research. The results for the much longer 460B run, where performance also matches, and detailed benchmark numbers are provided in Appx. B.5.
8B. While a full 8B training run is beyond our limits, we provide a first investigation in Fig. 9: In line with all our experiments, we see matching loss values and, promisingly, no instability at such a large scale, though the two runs are clearly short (20k steps). Moreover, instabilities arising from a high LR for longer training can be alleviated by methods like QK norm (Wortsman et al., 2024).
@@ 4 Do We Even Need to Cooldown?
In Section 3, we show that a constant LR with a short cooldown phase can replace the cosine decay. Ideally, however, we would not need a decay at all, but aim to obtain an optimal model at any point in training, even when prolonged. This would save even more computational resources and time. In this section, we investigate two potential approaches: weight averaging and a schedule-free optimizer.
@@ 4.1 Stochastic Weight Averaging (SWA)
Motivation. While slow annealing of the learning rate can be essential to find good minima (Smith et al., 2017; Loshchilov & Hutter, 2016), Sandler et al. (2023) show theoretical and empirical equivalence between stochastic weight averaging (SWA) and cooldowns in training vision models. There, intuitively and naturally averaging reduces noise and thus improves generalization (Wortsman et al., 2022). Motivated by these results, we aim to answer the same question in LLM training: Can weight averaging replace the cooldown phase?
Method. We opt for a form of SWA (Izmailov et al., 2018) that splits the training in fixed windows and averages within a window, allowing us to keep the average as a single additional copy of the model parameters. In our experiments, we set the window to $h=500$ steps and save the window averages as checkpoints every $h$ steps, which allows ad hoc evaluation of longer windows to see potential improvements akin to latest weight averaging (LAWA, Kaddour, 2022; Sanyal et al., 2023). For all our experiments, we find that windows below or at 2500 steps (256M tokens) are optimal. We also experimented with an exponential moving average (EMA), which performed worse than SWA, and therefore we do not report EMA.
Experimental results. We evaluate SWA with the same 210M model and show the results in Fig. 10 for both a constant LR (left) and cosine (right). Notably, we find a significant performance boost for SWA on top of a constant LR. Yet, it does not reach the loss values of explicit cooldowns. On the other hand, in line with previous work (Kaddour, 2022; Sanyal et al., 2023; Andriushchenko et al., 2023), we see a similar boost for SWA on top of cosine. This suggests that, regardless of schedule, SWA is a compelling approach to achieving strong models along the data-scale axis that can serve as a replacement for models trained with fewer steps, if the performance gap is acceptable and one wishes to avoid cooldowns. This is particularly advantageous as it can arguably be done for free on top of existing well-tuned optimizers.
@@ 4.2 Schedule-Free Optimizer (SFO)
Results. We compare the results of a long 210M model run with cooldown vs. SFO with AdamW in Figure 11. Although the optimizer does not require a learning rate schedule, the authors point out that training is more sensitive to the choice of the momentum parameters $(\beta_{1},\beta_{2})$ (which are not identical to the momentum in Adam). They note that the optimal parameters may depend on the length of training, making it not fully schedule-free. We observe this sensitivity in our experiments, where the choice of $(0.9,0.95)$ performs significantly worse and even increases the loss toward the end of training. For $(\beta_{1}=0.95,\beta_{2}=0.99)$, SFO performs remarkably well. Nevertheless, both settings are matched or outperformed by the cooldown schedule, in particular when comparing the same momentum configuration. We did not perform any further hyperparameter tuning for either method.
@@ 5 The Implications for Scaling Law Research
where $\{A,\alpha,B,\beta,E\}$ are variables to be estimated (Kaplan et al., 2020). Such scaling laws serve a multitude of critical purposes — from optimally spending a fixed amount of compute (FLOPs) for achieving the lowest loss, to trading off data-sources of different quality (Bi et al., 2024; Goyal et al., 2024) or modalities (Aghajanyan et al., 2023), to comparing the efficiency of different architectures (Gu & Dao, 2023; De et al., 2024).
Crucially, Hoffmann et al. (2022) demonstrated that it is necessary to vary the number of training tokens for a fixed family of models and fully decay the LR (see also Kaddour et al., 2023). At the time, their results suggested that LLMs were over-sized, leading to a substantial increase in data collection and training for much longer, beyond the Chinchilla optimal point $(N,D)$ (Touvron et al., 2023a, b).
@@
Experimental setup. We mimic a small-scale experimental setup for scaling laws: We train a range of model sizes (33M-360M) across different token scales (0.3B-10B) on the same SlimPajama 6B dataset. For each model, we choose exactly three token counts (around the Chinchilla optimal ratio of D/N=20) in increments of 10, 20 and 30 tokens per parameter. With the cosine schedule, each model is trained from scratch three times. In contrast, for averaging, we train each model just once for the longest cycle and then use the averages at the same token count as cosine; for cooldown, we similarly take checkpoints along the constant LR trajectory and perform three annealing periods to match the token counts. We adjust the LR to be higher for smaller models. In line with the findings of Sect. 3 and for a fair comparison, we set the constant LR to be half the maximum LR for cosine and perform 20% (linear) cooldown steps. More details are given in Appendix A.1.
Figure 12: Cooldown LR schedule + SWA scale reliably. We train a range of models (33M-360M), each for three different cosine lengths. We then run each model with a constant LR just once and take the snapshots at the same token count as cosine (for SWA) or perform post-train cooldowns to the same length. Each final model is represented by a dot. Left: The loss curve envelopes. Right: The perplexity of cosine (y-axis) vs. cooldowns and SWA. Points on the diagonal indicate the same loss for both methods; above outperforming cosine, below fall short. We see alignment between both methods.
@@
Additional results. We plot the training curves of all models in Appx. B.3. In addition, we show the same findings with experiments on OpenWebText2 in Appx. B.4.